OpenStudy (anonymous):

A rectangular piece of fabric is 16ft by 11ft. A triangular area with a height of 5.4ft and a base of 8.2ft is cut from the fabric. How much area is left over? A. 131.72 ft2 B. 153.86 ft2 C. 221.1 ft2 D. 263.44 ft2

OpenStudy (anonymous):

B. n ill explain ok area of a rectangle is A=L*W so A=16*11 n its 176 n for the triangle its A=1/2*B*H so A=0.5*5.4*8.2 n its 22.14 n if you subtract 176-22.14 you get 153.86